[figure: 9.7 cm]

VERTICAL AND DIAGONAL STATIC STRIPES IN THE HUBBARD MODEL

Marcin Raczkowski,^{a} Bruce Normand,^{b} and Andrzej M. Oleś^{a} aInstitute of Physics, Jagellonian University, Reymonta 4, PL-30059 Kraków, Poland

bInstitut de Physique Théorique, Pérolles, CH-1700 Fribourg, Switzerland

Qualitative picture of stable static stripe phases which occur in doped superconducting cuprates can be given within the Hubbard model using Hartree-Fock approximation [1]. Here we extend earlier Hartree-Fock studies of the stripe stability [2], and determine the phase diagram of the Hubbard model with nearest-neighbor ∝ t, by varying the on-site Coulomb repulsion U and next-neighbor ∝ t′ hopping, by investigating the locally stable structures for a few representative doping levels. Qualitatively vertical (01) [diagonal (11)] stripes are stable for U < 4t [U > 4t], if t′ = 0. Negative t′ (t′/t < 0), as obtained for the realistic parameters of high temperature superconductors, stabilizes vertical (01) stripes. In the anisotropic Hubbard model (t_{x} = 1.1t, t_{y} = 0.9t) vertical (01) stripes are more stable than diagonal stripes even at t′ = 0. Furthermore, we report the changes in the stability of (11) and (01) stripes due to repulsive (V > 0) and attractive (V < 0) nearest-neighbor interaction. If V > 0 the diagonal stripes are stable, while vertical (01) stripes are stabilized by the weak attractive interaction V ≃ −0.5t.

[1] A.M. Oleś, Acta Phys. Polon. B 31, 2963 (2000).

[2] B. Normand and A.P. Kampf, Phys. Rev. B 64, 024521 (2001).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Marcin Raczkowski

Address for correspondence : Institute of Physics Jagellonian University Reymonta 4 PL-30059 Kraków

Email address : marcin@alphetna.if.uj.edu.pl